OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to3x+2y=6 and passes through the point (3, 0).

OpenStudy (anonymous):

x/3/4 -y/2=1

OpenStudy (aroub):

2y=-3x+6 y=-3/2 +3 y-y1=m(x-x1), m= 2/3 , ( 3,0) y-0=2/3(x-3)